# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

### v2

@@ Abstract
/Understanding camera dynamics is a fundamental pillar of video spatial intelligence. However, existing multimodal models predominantly treat this task as a black-box classification, often confusing physically distinct motions by relying on superficial visual patterns rather than geometric cues. We present CamReasoner, a framework that reformulates camera movement understanding as a structured inference process to bridge the gap between perception and cinematic logic. Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to decode spatio-temporal cues such as trajectories and view frustums within an explicit reasoning block. To instill this capability, we construct a Large-scale Inference Trajectory Suite comprising 18k SFT reasoning chains and 38k RL feedback samples. Notably, we are the first to employ RL for logical alignment in this domain, ensuring motion inferences are grounded in physical geometry rather than contextual guesswork. By applying Reinforcement Learning to the Observation-Think-Answer (O-T-A) reasoning paradigm, CamReasoner effectively suppresses hallucinations and achieves state-of-the-art performance across multiple benchmarks.
@@ 1 Introduction
Camera movement serves as the invisible brushstroke of cinematic storytelling, acting as a vital conduit to convey subtle emotions and orchestrate audience immersion. Whether through Scorsese’s seamless tracking shots that create a seductive flow, or Spielberg’s dynamic dollies that heighten suspense and awe, these trajectories represent the intentionality of a director shaping the viewer’s perception of the physical and emotional world. In the landscape of Multimodal Large Language Models (MLLMs), camera movement understanding has emerged as a critical frontier that bridges the gap between passive pixel perception and active semantic reasoning. From the perspective of understanding, mastering these dynamics allows MLLMs to effectively decouple complex object movements from the observer’s ego-motion, even when camera intrinsic parameters are unknown, thereby enabling a more robust comprehension of 3D spatial geometry and narrative intent. This transition from surface-level recognition to deep structural analysis transforms an AI from a mere observer into a cinematic reasoner capable of interpreting the "why" behind every frame. Furthermore, from the perspective of generation, camera movement acts as the essential cornerstone for controllable video synthesis. As the frontier of AI shifts toward purposeful visual artistry, the ability to generate physically consistent videos relies fundamentally on the model’s grasp of how trajectories interact with scene depth and perspective. By integrating these motion priors, MLLMs can move beyond disorganized pixel synthesis toward achieving true world modeling capabilities, where every shift in perspective is both physically grounded and narratively coherent.
@@
Despite the growing demand for such sophisticated motion perception, current methodologies for camera movement understanding remain split between two constrained paradigms that struggle to meet these requirements. Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm, zhang2023protein] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy. These systems are fragile in dynamic environments because they often fail to disentangle the camera’s ego-motion from the complex movements of subjects within the scene. In contrast, modern learning-based Video-Language Models (VLMs) [comanici2025gemini, qwen3vl, hurst2024gpt] offer a promising alternative by treating motion understanding as a semantic task. Leveraging vast pre-training corpora, these models achieve impressive alignment between visual cues and cinematic terminology, while even demonstrating an emerging capacity for high-level reasoning [liu2025shotbench]. However, these models predominantly operate as black boxes that leap from pixels to labels without explicit reasoning. This shortcut often triggers hallucinations because the model relies on contextual cues like a running athlete instead of decoding the actual physics of the lens. Consequently, they struggle to distinguish physically distinct yet visually similar operations such as a zoom versus a dolly. These limitations necessitate a framework that bridges the gap between raw perception and structured cinematic logic.
@@
To bridge this gap, we introduce CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. Central to our approach is the Observation-Thinking-Answer (O-T-A) paradigm, which we implement through a sequential training pipeline of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). To support this process, we constructed a Large-scale Inference Trajectory Suite comprising 18k SFT and 38k RL samples. While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ground final verdicts in 3D spatio-temporal physics. By explicitly tracing trajectories and analyzing view frustums, our model establishes a rigorous logical premise before delivering a verdict, effectively eliminating contextual guesswork. As the first to apply Reinforcement Learning for logical alignment in this domain, CamReasoner-7B achieves state-of-the-art results, reaching 78.4% in binary classification and 74.5% in VQA. Notably, it excels in challenging Confusable Motion scenarios (60.7%) by distinguishing physical movements like truck versus pan through geometric reasoning. By integrating structural motion priors, CamReasoner provides a scalable path for MLLMs to achieve precise cinematic control and physically-consistent world modeling.
@@
  • RL-driven Logical Alignment: We are the first to employ Reinforcement Learning for logical alignment in this domain, achieving state-of-the-art 78.4% and 74.5% accuracy in binary classification and visual question answering tasks.
@@ 2.1 Camera Movement Understanding
Camera movement understanding bridges geometric perception and semantic reasoning, yet current methods face significant bottlenecks. Traditional geometric approaches rely on SfM-based pseudo ground-truth, which is largely restricted to static scenes and ignores scene context [zhou2018stereo, ling2024dl3dv, jin2024stereo4d, schonberger2016structure]. Categorical recognition tasks, while providing semantic labels, often suffer from poor specification—confusing rotation with translation (e.g., pan vs. truck) and failing to capture co-occurring motions in complex cinematography [huang2020movienet, rao2020shot, argaw2022anatomy]. Although recent benchmarks like ShotBench [liu2025shotbench, wu2025refineshot] evaluate nuanced cinematic grammar, existing methodologies lack a structured reasoning process to explain the visual cues driving their interpretations [hong2025motion, wang2024tarsier]. Consequently, there remains a critical need for a framework that integrates spatial evidence with logical deduction to achieve a holistic understanding across diverse video domains [zhou2018stereo, huang2020movienet, argaw2022anatomy, liu2025shotbench].
@@ 2.2 Multimodal Reinforcement Learning
Inspired by the reasoning success of LLMs, recent works have extended structured thinking to MLLMs across diverse visual tasks [li2025star, sun2025reinforcement, feng2025video, sun2025spacevista, zhou2025reinforced, duan2025codeplot, chen2025advancing, chen2025ares, meng2025open]. These include task-specific reasoning models for image VQA [huang2025vision, mei2025surveycontextengineeringlarge, mei2025a1], video understanding [feng2025video, li2025videochat], object detection [yu2025perception, shen2025vlm], segmentation [you2025seg], and temporal grounding [wang2025time]. Notably, frameworks like SophiaVL-R1 [fan2025sophiavl], OneThinker [feng2025onethinker], and FrameMind [ge2025framemind] introduce unified reasoning via reinforcement learning, while Thinking-with-Sound [xiong2025thinking] extends this paradigm to the audio domain through acoustic tool manipulation. However, despite the surge in RL-based multimodal reasoning, its application to the specialized domain of camera movement understanding remains unexplored.
@@ 3 Method
We introduce a two-stage training strategy combining an SFT cold start with Group Relative Policy Optimization (GRPO) [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding. This pipeline ensures the model efficiently acquires structured reasoning traces while maintaining format adherence and task accuracy. To support this, we curate distinct datasets specifically tailored for sequential Supervised Fine-Tuning and Reinforcement Learning phases.
@@ 3.1 Data Collection and Curation
3.1.1 CamReasoning-SFT-18k Construction.
To instill structured reasoning, we curate the CamReasoning-SFT-18k dataset via an answer-conditioned synthesis pipeline. As shown in Fig. 2, we utilize Qwen2.5-VL-72B to generate 38,672 initial trajectories from original video-QA pairs, enforcing the <observation>, <think>, and <answer> format. To ensure logical coherence, we implement a multi-dimensional filtration process using Qwen3 [qwen3vl] to verify samples based on format adherence, label accuracy, and reasoning consistency. This results in 18,541 high-fidelity samples that serve as cold-start data for teaching the model step-by-step cinematic inference.
3.1.2 CamReasoning-RL-38k
For the reinforcement learning stage, we construct the CamReasoning-RL-38k dataset to drive GRPO [shao2024deepseek]. Derived from the CameraBench training split, this dataset provides a diverse array of camera movement scenarios converted into our standardized instruction schema. Unlike the SFT phase, these samples lack reasoning annotations; instead, they serve as environmental queries for the model to refine its self-generated reasoning paths. This setup allows the model to optimize logical robustness and prediction accuracy through trial and error, guided by task-specific reward signals.
@@ 3.2.2 Reinforcement Learning Framework
To further align the model’s reasoning capabilities with task requirements and ensure robust logical chains, we employ RL. The training framework is structured into the following components:
Problem Formulation.
We formulate the camera movement understanding task as a sequential decision-making process. Given an input video clip $v$ and a textual query $q$ regarding the camera motion, the multimodal policy $\pi_{\theta}$ generates a structured response sequence $o$. This output $o$ comprises the visual observation, the explicit reasoning trace, and the final predicted camera movement category. Our optimization objective is to maximize the expected reward of these generated sequences under the policy:
@@
Reward Function Design.
Designing an effective reward function is critical for guiding the O-T-A reasoning process. We define a composite reward $R(o)$ that evaluates both the structural validity and the factual correctness of the output. The reward consists of two components: (1) Format Reward $r_{fmt}$: To enforce the strict reasoning structure, we apply a regular expression check. The model receives a reward of $1.0$ if and only if the generated output strictly adheres to the sequence of <observation>, <think>, and <answer> tags; otherwise, $r_{fmt}=0.0$. (2) Accuracy Reward $r_{acc}$: Conditional on passing the format check, we extract the prediction enclosed within the <answer> tags. If the predicted content matches the ground truth camera movement label, $r_{acc}$ is set to a positive score (e.g., $1.0$); otherwise, a penalty is applied. If the format check fails, $r_{acc}$ is explicitly set to $0.0$.
The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight. The overall reward is formulated as:
$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}.$ (3)
This formulation balances the trade-off between structure adherence and prediction accuracy, ensuring the model learns to generate valid reasoning traces while optimizing for correct answers.
GRPO.
To optimize the policy efficiently without the memory overhead of a value network, we utilize the GRPO algorithm. For each query $q$, we sample a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\theta_{old}}$. GRPO estimates the baseline using the group average, avoiding the need for a separate critic model. The training objective is to minimize the following surrogate loss: Let $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ be the probability ratio.
$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},$ (4)
$\quad\text{clip}\big(r_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$
where $\epsilon$ is the clipping parameter, $A_{i}$ is the advantage value, and $\beta$ controls the strength of the KL divergence penalty $D_{KL}$ (measured against the reference model $\pi_{ref}$).
Stability Strategy: EMA-GRPO.
To ensure training stability, we adopt the EMA-GRPO [feng2025onethinker] strategy for advantage normalization. Standard GRPO normalizes rewards using the immediate batch standard deviation. This introduces significant gradient fluctuations, as low-variance rollouts can lead to artificially inflated advantages and unstable updates. To mitigate this, we employ an exponential moving average (EMA) to estimate a globally stable standard deviation, denoted as $\sigma(t)$. The advantage is computed as:
$A_{i}(t)=\frac{R_{i}-\mathrm{mean}(\{R_{j}\})}{\sigma(t)}.$ (5)
Here, $\sigma(t)$ captures the intrinsic reward scale and filters out high-frequency noise from individual batches. EMA-GRPO effectively enforces a globally consistent trust region by stabilizing the reward scale, preventing spurious large policy updates caused by accidental low-entropy rollout groups. This approach decouples gradient scaling from immediate rollout randomness and ensures robust convergence during single-task training.
@@ 4 Experiments
4.1 Setup
Evaluation Settings
We evaluate our model using data from the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks. We tested the accuracy of multiple existing MLLMs using the questions and videos across the datasets. To ensure visual grounding, each question is paired with two videos yielding opposite answers. During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128. This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.

### v3

@@ Abstract
/Understanding camera dynamics is a fundamental pillar of video spatial intelligence. However, existing multimodal models predominantly treat this task as a black-box classification, often confusing physically distinct motions by relying on superficial visual patterns rather than geometric cues. We present CamReasoner, a framework that reformulates camera movement understanding as a structured inference process to bridge the gap between perception and cinematic logic. Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to articulate spatio-temporal observations and reason about motion patterns within an explicit reasoning block. To instill this capability, we construct a Large-scale Inference Trajectory Suite comprising 18k SFT reasoning chains and 38k RL feedback samples. To the best of our knowledge, we are the first to employ RL for logical alignment in camera movement understanding, ensuring motion inferences are grounded in structured visual reasoning rather than contextual guesswork. Built upon Qwen2.5-VL-7B, CamReasoner-7B improves binary classification accuracy from 73.8% to 78.4% and VQA accuracy from 60.9% to 74.5% over its backbone, consistently outperforming both proprietary and open-source baselines across multiple benchmarks.
@@ 1 Introduction
Camera movement serves as the invisible brushstroke of cinematic storytelling, acting as a vital conduit to convey subtle emotions and orchestrate audience immersion. Whether through Scorsese’s seamless tracking shots that create a seductive flow, or Spielberg’s dynamic dollies that heighten suspense and awe, these trajectories represent the intentionality of a director shaping the viewer’s perception of the physical and emotional world. In the landscape of Multimodal Large Language Models (MLLMs), camera movement understanding has emerged as a critical frontier that bridges the gap between passive pixel perception and active semantic reasoning. From the perspective of understanding, mastering these dynamics allows MLLMs to effectively decouple complex object movements from the observer’s ego-motion, even when camera intrinsic parameters are unknown, thereby enabling a more robust comprehension of 3D spatial geometry and narrative intent. This transition from surface-level recognition to deep structural analysis moves an AI beyond surface-level pattern matching toward more structured motion analysis. Furthermore, from the perspective of generation, camera movement acts as the essential cornerstone for controllable video synthesis. As the frontier of AI shifts toward purposeful visual artistry, the ability to generate physically consistent videos relies fundamentally on the model’s grasp of how trajectories interact with scene depth and perspective. By integrating these motion priors, MLLMs can move beyond disorganized pixel synthesis toward achieving true world modeling capabilities, where every shift in perspective is both physically grounded and narratively coherent.
@@
Despite the growing demand for such sophisticated motion perception, current methodologies for camera movement understanding remain split between two constrained paradigms that struggle to meet these requirements. Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy. These systems are fragile in dynamic environments because they often fail to disentangle the camera’s ego-motion from the complex movements of subjects within the scene. In contrast, modern learning-based Video-Language Models (VLMs) [comanici2025gemini, qwen3vl, hurst2024gpt] offer a promising alternative by treating motion understanding as a semantic task. Leveraging vast pre-training corpora, these models achieve impressive alignment between visual cues and cinematic terminology, while even demonstrating an emerging capacity for high-level reasoning [liu2025shotbench]. However, these models predominantly operate as black boxes that leap from pixels to labels without explicit reasoning. This shortcut often triggers hallucinations because the model relies on contextual cues like a running athlete instead of decoding the actual physics of the lens. Consequently, they struggle to distinguish physically distinct yet visually similar operations such as a zoom versus a dolly. These limitations necessitate a framework that bridges the gap between raw perception and structured cinematic logic.
@@
To bridge this gap, we introduce CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. Central to our approach is the Observation-Thinking-Answer (O-T-A) paradigm, which we implement through a sequential training pipeline of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). To support this process, we constructed a Large-scale Inference Trajectory Suite comprising 18k SFT and 38k RL samples. While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ensure final verdicts are grounded in structured visual reasoning. By articulating detailed visual observations and reasoning about motion patterns, our model establishes a rigorous logical premise before delivering a verdict, substantially reducing reliance on contextual guesswork. To the best of our knowledge, we are the first to apply Reinforcement Learning for logical alignment in camera movement understanding. CamReasoner-7B achieves the best results among all compared methods, reaching 78.4% in binary classification and 74.5% in VQA. Notably, it excels in challenging Confusable Motion scenarios (60.7%) through structured spatial reasoning. By integrating structural motion priors, CamReasoner provides a scalable path for MLLMs to achieve precise cinematic control and physically-consistent world modeling.
@@
  • RL-driven Logical Alignment: To the best of our knowledge, we are the first to employ Reinforcement Learning for logical alignment in camera movement understanding, achieving 78.4% and 74.5% accuracy in binary classification and visual question answering under our reformulated evaluation protocol, outperforming all compared baselines.
@@ 2.1 Camera Movement Understanding
Camera movement understanding bridges geometric perception and semantic reasoning, yet current methods face significant bottlenecks. Traditional geometric approaches rely on SfM-based pseudo ground-truth, which is largely restricted to static scenes and ignores scene context [zhou2018stereo, ling2024dl3dv, jin2024stereo4d, schonberger2016structure]. Categorical recognition tasks, while providing semantic labels, often suffer from ambiguous taxonomy—confusing rotation with translation (e.g., pan vs. truck) and failing to capture co-occurring motions in complex cinematography [huang2020movienet, rao2020shot, argaw2022anatomy]. Recent studies further expose these weaknesses: Feng et al. [feng2026geometry] demonstrate that VideoLLM vision encoders weakly encode camera motion cues and propose injecting geometric signals from 3D foundation models to compensate, while VEU-Bench [li2025veu] reveals that Vid-LLMs struggle with editing-level understanding including shot motion, often performing below chance. Although recent benchmarks like ShotBench [liu2025shotbench, wu2025refineshot] evaluate nuanced cinematic grammar, existing methodologies lack a structured reasoning process to explain the visual cues driving their interpretations [hong2025motion, wang2024tarsier]. Consequently, there remains a critical need for a framework that integrates spatial evidence with logical deduction to achieve a holistic understanding across diverse video domains [zhou2018stereo, huang2020movienet, argaw2022anatomy, liu2025shotbench].
@@ 2.2 Multimodal Reinforcement Learning
Inspired by the reasoning success of LLMs, recent works have extended structured thinking to MLLMs across diverse visual tasks [li2025star, sun2025reinforcement, feng2025video, sun2025spacevista, zhou2025reinforced, duan2025codeplot, chen2025advancing, chen2025ares, meng2025open]. These include task-specific reasoning models for image VQA [huang2025vision, mei2025surveycontextengineeringlarge, mei2025a1], video understanding [feng2025video, li2025videochat], object detection [yu2025perception, shen2025vlm], segmentation [you2025seg], and temporal grounding [wang2025time]. Notably, frameworks like SophiaVL-R1 [fan2025sophiavl], OneThinker [feng2025onethinker], and FrameMind [ge2025framemind] introduce unified reasoning via reinforcement learning, while Thinking-with-Sound [xiong2025thinking] extends this paradigm to the audio domain through acoustic tool manipulation. However, despite the surge in RL-based multimodal reasoning, its application to camera movement understanding remains largely underexplored.
@@ 3 Method
We introduce a two-stage training strategy combining an SFT cold start with GRPO [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding. This pipeline ensures the model efficiently acquires structured reasoning traces while maintaining format adherence and task accuracy. To support this, we curate distinct datasets specifically tailored for sequential training.
@@ 3.1 Data Collection and Curation
CamReasoning-SFT-18k.
To instill structured reasoning, we curate the CamReasoning-SFT-18k dataset via an answer-conditioned synthesis pipeline. As shown in Fig. 2, we utilize Qwen2.5-VL-72B to generate 38,672 initial trajectories from original video-QA pairs, enforcing the <observation>, <think>, and <answer> format. To ensure logical coherence, we implement a multi-dimensional filtration process using Qwen3 [qwen3vl] to verify samples based on format adherence, label accuracy, and reasoning consistency. We further conduct human verification on a random subset of 500 samples, achieving an acceptance rate of 93%, confirming the reliability of the automatic filtration. This results in 18,541 high-fidelity samples that serve as cold-start data for teaching the model step-by-step cinematic inference.
CamReasoning-RL-38k
For the reinforcement learning stage, we construct the CamReasoning-RL-38k dataset to drive GRPO [shao2024deepseek]. Derived from the same CameraBench training split used for SFT data generation, this dataset comprises 38,451 video-QA pairs converted into our standardized instruction schema. Unlike the SFT phase, these samples lack reasoning annotations; instead, they serve as environmental queries for the model to refine its self-generated reasoning paths. This setup allows the model to optimize logical robustness and prediction accuracy through trial and error, guided by task-specific reward signals.
@@ 3.2.2 Reinforcement Learning Framework
Building upon the SFT-initialized policy, we employ RL to further align the model’s reasoning with task requirements. Given an input video clip $v$ and a textual query $q$, the policy $\pi_{\theta}$ generates a structured response $o$ comprising observation, reasoning trace, and predicted camera movement. The optimization objective is:
@@
Reward Design.
We define a composite reward with two components: (1) a Format Reward $r_{fmt}$ that equals $1.0$ when the output strictly follows the <observation>-<think>-<answer> structure and $0.0$ otherwise, and (2) an Accuracy Reward $r_{acc}$ that evaluates whether the prediction within <answer> matches the ground truth, conditional on passing the format check. The final reward is:
$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt},$ (3)
where $\lambda$ balances structure adherence and prediction accuracy.
Optimization.
We adopt GRPO to optimize the policy without the memory overhead of a value network. For each query, we sample a group of $G$ outputs from the current policy and estimate the baseline using the group average. The surrogate loss is:
$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},\text{clip}\big(r_{i}(\theta),1\!-\!\epsilon,1\!+\!\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$ (4)
where $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ is the probability ratio, $\epsilon$ is the clipping parameter, and $\beta$ controls the KL penalty against the reference policy. To ensure training stability, we adopt EMA-GRPO [feng2025onethinker], which replaces the batch-level standard deviation with an exponential moving average estimate $\sigma(t)$ for advantage normalization: $A_{i}(t)=({R_{i}-\mathrm{mean}(\{R_{j}\})})/{\sigma(t)}$. This stabilizes gradient scaling and prevents spurious policy updates caused by low-variance rollouts.
@@ 4 Experiments
4.1 Experimental Setup
Benchmark and Dataset
We evaluate on three benchmarks. CameraBench [lin2025towards] covers two tasks: 1) binary classification, where the model determines whether a specific motion primitive such as “pan left” is present in a given video, evaluated on 7,679 test samples; and 2) multi-choice VQA, where the model selects the correct camera movement description from several candidates, evaluated on 14,172 test samples spanning 9 categories. We modify the original benchmark by replacing mAP with accuracy for binary classification and reformulating VQA from pairwise confidence-score comparison into a standard multiple-choice format. These changes shift the evaluation focus from score calibration to decision-making correctness, enabling a more direct measure of whether a model can correctly interpret camera movements. For both tasks, we report the overall average accuracy weighted by the number of test samples in each category. Further details are provided in Appendix B. All baseline results are obtained by re-running each model under our reformulated protocol with identical prompts and decoding settings. ShotBench [liu2025shotbench] is a benchmark for cinematic shot understanding; we evaluate on its camera movement subset using accuracy. RefineShot [wu2025refineshot] extends ShotBench to assess reasoning reliability: an external LLM judges whether the model produces consistent reasoning across semantically equivalent queries, and the reliability score is the proportion of consistent outputs among all evaluated samples.